OpenStudy (anonymous):

simplify completely 3x^2+21x-54/x^2+3x-54

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{3x^2+21x-54}{x^2+3x-54 } } \] \[\huge\color{blue}{ \frac{3(x^2+7x-18)}{x^2+3x-54 } } \] \[\huge\color{blue}{ \frac{3( x+9 ) (x-2)}{(x-6)(x+9) } } \]

OpenStudy (solomonzelman):

Go from here.

OpenStudy (shamil98):

was gonna say let her factor it out, she's asked previous questions with the same thing, ...

OpenStudy (solomonzelman):

Oh, I didn't see the previous question, just logged on, sorry.

OpenStudy (solomonzelman):

@helpineedit, can you do it from where I left of?

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{3 \color{red} { (x+9) } (x-2)}{(x-6)\color{red} { (x+9) } } } \]see what cancels?

OpenStudy (anonymous):

so 3(x-2)/(x-6)

OpenStudy (solomonzelman):

Yup!

OpenStudy (anonymous):

thnks for helping me :)

OpenStudy (solomonzelman):

YW!

OpenStudy (anonymous):

could u help me with 3 more?

OpenStudy (solomonzelman):

i can try at least.

OpenStudy (solomonzelman):

Ask

OpenStudy (anonymous):

simplify completely 2x^2+13x+20/x^2-5x-36

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{2x^2+13x+20}{x^2-5x-36 } } \]can you do this the same way I did the previous question?

OpenStudy (anonymous):

2+13x+20/-5x-36

OpenStudy (solomonzelman):

What? Are you saying that that's the question (instead of what I wrote) ?

OpenStudy (anonymous):

oh no i was trying and i guess failing at canceling out

OpenStudy (solomonzelman):

Sorry, hold on.

OpenStudy (shamil98):

Factor the numerator and denominator..

OpenStudy (solomonzelman):

\[\huge\color{red}{ \frac{2x^2+13x+20}{x^2-5x-36} } \] \[\huge\color{blue}{ \frac{ (2x~+~\text{_})(x~+~\text{_}) }{ (x~-~~\text{_})(x~+~\text{_})} }\]

OpenStudy (solomonzelman):

Fill in blanks.

OpenStudy (anonymous):

(2x+13x)(x+20) (x-5x)(x+36)

OpenStudy (solomonzelman):

Wow, no, you are incorrect.

OpenStudy (anonymous):

darn! it was kind of confusing what number to put were sorry

OpenStudy (solomonzelman):

Lets try to think. For the top 2a+b=13 a*b=20 For the bottom a-b=5 a*b=36

OpenStudy (solomonzelman):

(a and b are the 2 constants)

OpenStudy (anonymous):

ok i tried to do the problem and this is what i got 2(x+10)/x-9

OpenStudy (solomonzelman):

No.

OpenStudy (solomonzelman):

Can you factor it out? (Again) For the top 2a+b=13 a*b=20 For the bottom a-b=5 a*b=36

OpenStudy (anonymous):

ohh yea your right.

OpenStudy (anonymous):

is there a 13 in it at all?

OpenStudy (solomonzelman):

What do you mean?

OpenStudy (anonymous):

like im not sure if its 13x+5 or 2(13x+5)

OpenStudy (solomonzelman):

Do you know how to factor at all? To me you seem unprepared for this, do you want to learn how to factor?

OpenStudy (anonymous):

yes please

OpenStudy (solomonzelman):

Hold on, let's see, maybe http://www.youtube.com/watch?v=v0dhpNnzZjA

OpenStudy (solomonzelman):

Or I can make a question where you would get an IS tutor, I am not good at teaching from zero, sorry.

OpenStudy (anonymous):

its ok i have to leave in a bit and need this done so maybe i can find another way

OpenStudy (solomonzelman):

Alright.